Shift report 20000714 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Fri Jul 14 2000 - 15:48:02 EDT

  • Next message: http account: "Shift report 20000715 16:00-24:00"

    ------------------------------------------------------------
    Date: Fri Jul 14 15:48:02 2000, Shift: 08:00-16:00
    Supervisor: Bjorn Samset, with: TV, PC et al.
    Log book#3 pages 1 - 7
    
    Shift Summary:
    ------------------------------------------------------------
    No beam all shift, except a few short injections.
    No accelerations.
    
    The broken TDC for the TOFW was fixed by KO, but due to
    lack of beam it is not yet really tested. PC worked on 
    debugging H1, but also lacked beam to complete tests.
    
    FV looked at run 1767 from last shift, and found clear 
    tracks in T2! A few pictures have been placed on the pii3 
    homepage.
    
    I have written a short instruction on how to start up
    HV-monitors on the tek17 xterminal. It is in logbook 3 on
    page 6, and on the pii3 homepage. I started a new section
    "additions" under the UserGuide as a place-holder for
    new writeups.
    
    Some engineers worked on the magnet power-supplies - no 
    problems reported.
    
    See detector-problems reported by MM last shift (logbook 2,
    inside of back cover...)
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Fri Jul 14 2000 - 15:51:52 EDT